A beaker contains 20mL of a solution that is 12% cetyl alcohol, a moisturizer often found in shampoo. How much of a solution that is 45% CEA should be added, so that the final mixture is 30% CEA?
Answer by stanbon(75887)

A beaker contains 20mL of a solution that is 12% cetyl alcohol, a moisturizer often found in shampoo. How much of a solution that is 45% CEA should be added, so that the final mixture is 30% CEA?
-----
Equation:
CEA + CEA = CEA
0.12*20 + 0.45x = 0.30(20+x)
----
12*20 + 45x = 30*20 + 30x
18*20 = 15x
x = 24 mL (amt of 45% solution needed)
========================================
Cheers,
Stan H.